A 2.50 kg block slides down an inclined plane at a 20.0 degree angle (constant velocity).

a) If the length of the incline is 2.0 m, calculate the work done by the force of gravity.

b) Calculate the work done by the force of friction.

c) Calculate the net work done by both gravity and friction.
Answer #1

(a)


along the incline component of gravitationl force Fg = m*g*sintheta


work done Wg = Fg*L = m*g*sintheta*L

Wg = 2.5*9.8*sin20*2 = 16.76 J


-----------


(b)

since the velocity is same


kinetic energy is same


Wtot = dKE

Wg + Wf = 0


Wf = -Wg = -16.76 J

(c)


Wnet = 0